Sixx AM Manufacturing has a target debt—equity ratio of 0.66. Its cost of equity is 18 percent, and its cost of debt is 10 percent. If the tax rate is 33 percent, what is the company's WACC?
Answer #1

After-tax cost of debt=10*(1-tax rate)

=10*(1-0.33)=6.7%

Debt-equity ratio=debt/equity

Hence debt=0.66 equity

Let equity be $x

Debt=$0.66x

Total=$1.66x

WACC=Respective cost*Respective weight

=(x/1.66x*18)+(0.66x/1.66x*6.7)

which is equal to

=13.51%(Approx).